OpenStudy (anonymous):

Find x correct to one decimal place. y = 60

OpenStudy (anonymous):

check attachment for triangle

OpenStudy (jack1):

use SOH CAH TOA sin (angle) = opposite side / hypotenuse so for left triangle : sin 60 = y / h sin 60 = 60 / h re-arrange the equation to work out h This will give you the length of one of the sides on the triangle on the right repeat for 2nd triangle

OpenStudy (jack1):

dude how'd you go?